# Energy-Efficient Balanced Flow Control Achieved through Optimization of Synthetic Jet Placement Based on Deep Reinforcement Learning

### v1

@@ 2.2 Circular cylinder environment
Figure 2: For a single pair of jets, five different arrangements are considered: (a) $J_{1}^{1}$ and $J_{1}^{2}$, (b) $J_{2}^{1}$ and $J_{2}^{2}$, (c) $J_{3}^{1}$ and $J_{3}^{2}$, (d) $J_{4}^{1}$ and $J_{4}^{2}$, (e) $J_{5}^{1}$ and $J_{5}^{2}$. For two pairs of jets, two arrangements are examined: (f) $J_{3}^{1}$, $J_{3}^{2}$, $J_{4}^{1}$, and $J_{4}^{2}$; (g) $J_{3}^{1}$, $J_{3}^{2}$, $J_{2}^{1}$, and $J_{2}^{2}$. For three pairs of jets, one arrangement is studied: (h) $J_{2}^{1}$, $J_{2}^{2}$, $J_{3}^{1}$, $J_{3}^{2}$, $J_{4}^{1}$, and $J_{4}^{2}$.
To thoroughly investigate the impact of synthetic jets positioning on flow control performance around a circular cylinder, a series of experiments is conducted, employing various jet configurations: one pair, two pairs, and three pairs of jets. In the single-pair configuration, five different jet placement schemes are examined to evaluate their effectiveness in controlling the flow. Each scheme is designated according to the position of the jets. For example, $J_{1}^{1}$ and $J_{1}^{2}$ refer to a pair of synthetic jets symmetrically positioned about the $x$-axis, with a net mass flow rate of zero across the pair. In the notation $J_{1}^{1}$, the subscript 1 indicates the first pair of synthetic jets, while the superscript 1 designates the first jet within that pair.
For each jet placement scheme, the azimuthal angles of the first jet in each pair are indicated in the figures. The configurations considered in this study include jets positioned at angles of 60°, 75°, 90°, 105°, and 120°. For the two-pair configuration, we explore two different placement schemes to evaluate the interaction effects between multiple pairs of jets. The three-pair configuration involves a specific arrangement designed to provide deeper insights into the cumulative effects of additional jet pairs on flow control. In this configuration, flow control is achieved by simultaneously activating $J_{2}^{1}$, $J_{2}^{2}$, $J_{3}^{1}$, $J_{3}^{2}$, $J_{4}^{1}$, and $J_{4}^{2}$. The precise locations and nomenclature of these jets are detailed in figure 2.
In scenarios involving multiple jet pairs, it is important to emphasize that the control strategy derived from DRL assigns distinct actions to each individual pair of jets. This approach allows the agent to assign distinct action values to jets positioned at various locations, facilitating tailored control that considers the unique dynamics associated with each jet placement. By assigning unique actions to each jet pair, the DRL framework facilitates the evaluation of the relative effectiveness of flow control at different positions. This method enables the quantification of the contribution of each jet pair to the overall control strategy, providing a deeper understanding of the role each jet plays in modulating the flow dynamics.
The objective of this study is to develop a DRL-based control strategy that enables simultaneous flow control at multiple positions on the cylinder. By assigning distinct actions to each individual jet pair, the agent can tailor the control to the specific dynamics associated with the placement of each jet. This approach allows for a systematic comparison of jet performance at various locations, facilitating the evaluation of the relative effectiveness of flow control across different positions. Furthermore, this comparison is complemented by an analysis of the underlying causes of observed performance differences, informed by principles of fluid mechanics. Such a comprehensive analysis enhances our understanding of how the contribution of each jet is influenced by its position and the physical mechanisms governing the flow dynamics.
@@ 2.3 Square cylinder environment
Figure 3: Schematic of the computational domain and boundary conditions. (a) Establishment of the coordinate system and detailed dimensions of the computational domain. (b) Placement of the synthetic jets near the trailing corner point. (c) The mesh generation scheme around the square object, illustrating the distribution of high-density mesh in regions prone to flow separation, and (d) A close-up view of the mesh structure, highlighting the refinement near the boundary layer.
The computational domain for simulating the flow around a square cylinder is precisely defined with dimensions of $40D\times 20D$, where $D$ represents the characteristic side length of the cylinder. The origin is located at the center of the cylinder, with $10D$ extending both above and below this point to define the vertical boundaries. The domain extends $10D$ upstream from the cylinder to the inlet boundary $\Gamma_{\text{in}}$, providing sufficient distance for the flow to fully develop into a uniform profile before interacting with the cylinder. Downstream, the domain stretches $30D$ from the cylinder to the outlet boundary $\Gamma_{\text{out}}$, ensuring ample space for capturing the wake formation and vortex shedding phenomena. As illustrated in figure 3(a), this spatial configuration is carefully optimized to minimize the influence of boundary effects on the flow around the cylinder while ensuring that the downstream region is sufficiently extended to capture the wake dynamics with high accuracy. This design is critical for maintaining the fidelity of the numerical simulations and ensuring the robustness and reliability of the results.
In the simulation setup, the inlet of the computational domain is prescribed with a uniform velocity profile, where the velocity magnitude is set to 2 and aligned along the $x$-axis. The synthetic jets are configured with velocities perpendicular to the walls of the square cylinder, directed along the outward normal to these walls, as illustrated in figure 3(b). Except at the locations of the synthetic jets, the remaining sections of the square cylinder walls are subjected to no-slip solid wall boundary conditions, ensuring that the fluid adheres to the surfaces and accurately reflecting the physical constraints imposed by solid boundaries on the fluid flow. At the outlet boundary, Neumann-type conditions are applied to maintain a zero stress vector, effectively simulating the natural flow behavior at an infinite distance. The upper and lower boundaries of the domain are defined with far-field boundary conditions, designed to minimize their impact on the flow around the square cylinder and closely approximate an unbounded fluid domain. This careful configuration ensures the accuracy and reliability of the simulation, allowing for a precise investigation of the flow dynamics.
@@
To ensure numerical stability, the time step size is carefully selected as $\Delta t=0.0005$ following extensive validation and iterative testing. This choice balances temporal accuracy with adherence to the CFL condition, particularly in regions of high flow velocity and steep gradients. The rigorous selection of this parameter minimizes numerical errors, thereby enhancing the fidelity of the simulation and ensuring a reliable analysis of the fluid dynamics. The mesh topology surrounding the square cylinder is illustrated in figure 3(c). The global mesh is designed to systematically coarsen with increasing radial distance from the cylinder, optimizing computational efficiency while maintaining accuracy. In contrast, figure 3(d) highlights the local mesh refinement near the cylinder, where the grid undergoes significant densification to form high-resolution elements. This local refinement is essential for accurately capturing the intricate flow dynamics within the boundary layer, particularly in areas with elevated shear and steep velocity gradients. Such precision in mesh design is crucial for resolving critical flow phenomena and ensuring the overall reliability of the simulation results.
To investigate the impact of synthetic jets positioning on the flow control performance around a square cylinder, we design a series of experiments with different jet configurations: one pair of jets, two pairs of jets, and three pairs of jets. In the one-pair jet configuration, we explore four different jet placement schemes. Each scheme involves a distinct spatial arrangement to assess the optimal positioning for effective flow control. The four different jet placement schemes are named according to the location of the synthetic jets. For instance, $J_{1}^{1}$ and $J_{1}^{2}$ represent a pair of jets located near the front corners of the square cylinder, where the subscript ”1” indicates the first pair of synthetic jets, which has a net mass flow rate of zero. The superscript ”1” or ”2” distinguishes between the two different jets within the pair.
For the two-pair jet configuration, we investigate a single jet placement scheme aimed at analyzing how two strategically positioned pairs of jets influence the flow dynamics. Similarly, for the three-pair jet configuration, we examine a specific jet placement scheme to understand the effects of three pairs of jets on flow control performance. The precise positions and designations of the jets for each configuration are detailed in figure 4. Notably, in scenarios involving multiple pairs of jets, the control strategy derived from DRL provides distinct control actions for each individual pair of jets. This approach allows for a nuanced assessment of how different configurations and placements of synthetic jets influence the overall control effectiveness in modifying the flow around the square cylinder.
@@ 2.4 Deep Reinforcement Learning
Deep learning is a critical branch of machine learning, with its core centered on utilizing the multi-layered structure of artificial neural networks to process and analyze data. This multi-layered structure, also known as deep neural network, consists of an input layer, multiple hidden layers, and an output layer (LeCun et al., 2015). The input layer receives raw data, such as image pixels or text word vectors, and introduces these features into the network. The hidden layers, positioned between the input and output layers, are composed of interconnected neurons and are responsible for the layer-by-layer processing and transmission of information. The output layer generates the final prediction results or classification labels, with its structure and the number of neurons tailored to the specific task (Granter et al., 2017). Through this layered processing and non-linear transformations, DNN can automatically extract complex features from raw data and capture intricate patterns, enabling deep learning to exhibit exceptional performance in tasks such as image recognition, natural language processing, and speech recognition.
Reinforcement Learning is a sophisticated machine learning paradigm that emphasizes empowering an agent to autonomously learn optimal actions within a dynamic environment through a process of trial and error (Mnih, 2013). The objective of the agent is to maximize cumulative rewards over time by continuously interacting with the environment, receiving feedback in the form of rewards or penalties, and incrementally refining the strategy to make increasingly effective decisions (François-Lavet et al., 2018; Wang et al., 2020). Through this iterative process, the agent adapts to the complexities of the environment and discovers the most advantageous actions to achieve long-term goals, ultimately enhancing its ability to make informed and efficient decisions across a wide range of scenarios.
@@
2.4.1 Foundation
The agent is the entity that executes actions, learning and optimizing its strategy through repeated interactions with the environment, as depicted in figure 5(a), which illustrates the basic structure of agent-environment interaction. The environment serves as the external system within which the agent operates, providing feedback in the form of rewards based on the actions taken by the agent ($a\in\mathbb{A}$) (François-Lavet et al., 2018). The state ($s\in\mathbb{S}$) represents a snapshot of the environment at a specific moment, typically perceived by the agent as information from the environment. The agent uses a policy ($\pi(a|s)$), which is a set of rules or a function that guides the selection of actions based on the current state. The policy can be either deterministic or stochastic. The state value function ($V(s)$) evaluates the desirability of a particular state, representing the expected cumulative reward that can be obtained starting from that state. Similarly, the action value function ($Q(s,a)$) assesses the desirability of taking a specific action in a given state, representing the expected cumulative reward that can be achieved by executing that action in the state (Wang et al., 2020).
2.4.2 Mathematical model
RL fundamentally relies on the concept of a finite markov decision process (MDP) to model the decision-making process (Puterman, 1990; Altman, 1999). A finite MDP is characterized by the tuple $(S,A,P,R,\gamma)$, where $S$ represents the finite state space, $A$ denotes the finite action space, $P$ is the state transition probability function, and $R$ is the reward function. The discount factor $\gamma$ quantifies the relative importance of future rewards compared to immediate rewards (Puterman, 1990). At each time step $i$, the agent observes the environment in state $s_{i}$, selects an action $a_{i}$, and transitions to the next state $s_{i+1}$ with probability $p(s_{i+1}|s_{i},a_{i})$. The agent then receives a reward $R(s_{i},a_{i},s_{i+1})$ associated with this transition. The decision-making process is governed by a policy $\pi(a_{i}|s_{i})$, which specifies the probability of choosing action $a_{i}$ given the current state $s_{i}$ (Altman, 1999).
2.4.3 Classification
RL can be broadly categorized into two main types: value-based methods and policy-based methods. Value-based methods involve learning a $V(s)$ or a $Q(s,a)$ that estimates the expected cumulative reward for a given state or state-action pair (Li, 2018; Altman, 1999). The agent selects the optimal action by maximizing these value functions, with examples including Q-learning and DQN (Mnih et al., 2015). Policy-based methods, on the other hand, directly learn a policy function ($\pi(a|s)$), which represents the probability distribution of selecting actions given a particular state (Li, 2018; Kaiser et al., 2024). The agent optimizes the policy function to maximize cumulative rewards, making this approach effective in handling continuous action spaces and stochastic policies. Common policy-based algorithms include policy gradient methods (Silver et al., 2014), trust region policy optimization (TRPO) (Schulman et al., 2015), and PPO (Schulman et al., 2017). The basic training structure of policy-based algorithms is illustrated in figure 5(b).
Figure 6: Basic structures of deep reinforcement learning.
DNNs are a class of machine learning models composed of multiple layers of interconnected neurons, capable of automatically learning and extracting features from data (LeCun et al., 2015). They excel in handling complex, high-dimensional datasets, making them essential for tasks such as image recognition, natural language processing, and reinforcement learning. In the framework of DRL, where DL plays a central role, the combined operation of DL and RL is depicted in figure 6. By using DNNs, deep learning can automatically extract key features from complex and high-dimensional data, significantly enhancing the ability of reinforcement learning to deal with complex environments. The functional representation ability of deep learning is extremely powerful, especially in policy representation and value function approximation, where it plays a crucial role. These capabilities enable DRL to achieve more accurate and efficient decision-making in high-dimensional, continuous action spaces, greatly improving its application effect and breadth in complex decision-making environments (Arulkumaran et al., 2017).
In reinforcement learning, the $V(s)$ and the $Q(s,a)$ are used to evaluate the goodness of a state or a state-action pair, typically represented in terms of cumulative rewards. The $V(s)$ represents the expected cumulative reward that can be obtained starting from state $s$ and following a policy $\pi$ (Szepesvári & Littman, 1999). It is defined as:
$V(s)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}\mid s_{0}=s\right],$ (10)
where $\mathbb{E}_{\pi}$ denotes the expectation under the policy $\pi$. $\gamma$ is the discount factor, $0\leq\gamma\leq 1$, which discounts future rewards. $r_{t}$ is the immediate reward received at time step $t$. $s_{0}=s$ indicates that the initial state is $s$.
The action value function $Q(s,a)$ represents the expected cumulative reward that can be obtained by taking action $a$ in state $s$ and subsequently following the policy $\pi$. It is defined as:
$Q(s,a)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}\mid s_{0}=s,a_{0}=a\right],$ (11)
@@ 2.4.3 Classification
where $\mathbb{E}_{\pi}$ similarly denotes the expectation under the policy $\pi$. $\gamma$ is the discount factor. $r_{t}$ is the immediate reward received at time step $t$. $s_{0}=s$ indicates that the initial state is $s$, and $a_{0}=a$ indicates that the initial action is $a$. These two functions together describe the expected cumulative rewards associated with different states and state-action pairs, helping the agent optimize its policy to maximize long-term returns.
The advantage function $A(s,a)$ is typically computed based on the $V(s)$ and the $Q(s,a)$. The advantage function $A(s,a)$ evaluates the relative benefit of taking action $a$ in a given state $s$ compared to other possible actions (Henderson et al., 2018). The standard formula is:
$A(s,a)=Q(s,a)-V(s),$ (12)
where $Q(s,a)$ is the action value function, representing the expected cumulative reward for taking action $a$ in state $s$. $V(s)$ is the state value function, representing the expected cumulative reward for being in state $s$ without considering which action is taken.
PPO is a powerful policy optimization algorithm that maintains the stability of policy performance by limiting the amplitude of policy updates. It uses a proxy objective function to encourage policy updates so that the probability ratio of the current policy to the old policy remains within a certain range, thereby avoiding drastic fluctuations in policy performance. Also, PPO uses the advantage function to evaluate the advantage of state-action pairs, and optimizes the proxy objective function through stochastic gradient descent, ultimately finding the optimal policy that can obtain the highest expected cumulative reward (Schulman et al., 2017).
$L^{\text{CLIP}}(\theta)=\hat{\mathbb{E}}_{t}\left[\min\left(\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}\hat{A}_{t},\text{clip}\left(\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})},1-\epsilon,1+\epsilon\right)\hat{A}_{t}\right)\right],$ (13)
where $L^{\text{CLIP}}(\theta)$ is used to optimize the performance of the current policy, where $\hat{\mathbb{E}}_{t}$ denotes the empirical expectation at time step $t$. The ratio of the probability of choosing action $a_{t}$ under the current policy $\pi_{\theta}(a_{t}|s_{t})$ to the probability under the old policy $\pi_{\theta_{\text{old}}}(a_{t}|s_{t})$ is represented as $\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}$, which measures the extent of policy change. The advantage function estimate $\hat{A}_{t}$ evaluates the relative value of taking action $a_{t}$ in state $s_{t}$. To maintain stability in policy updates, a clipping mechanism is introduced: $\text{clip}\left(\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})},1-\epsilon,1+\epsilon\right)$, which restricts the policy ratio within the range $[1-\epsilon,1+\epsilon]$, where $\epsilon$ is a small hyperparameter controlling the range of the update. This objective function allows the PPO algorithm to ensure stable policy updates while optimizing the policy performance.
2.5 DRL-Enhanced Active Flow Control
We couple a computational fluid dynamics environment with a DRL framework to address AFC. The two core components of DRL, the agent and the environment, are integral to this approach. In our study, we utilize a PPO agent to train the control policy, while the CFD environment simulates flow around either a circular or square cylinder. The primary task of the PPO agent is to determine an optimal real-time control strategy for two jet actuators positioned on the cylinder’s surface, with the goal of minimizing the fluid forces acting on the structure. The agent obtains state information by observing the CFD environment, with flow probes measuring quantities such as velocity and pressure within the flow field. The agent’s actions involve adjusting the mass flow rate of the synthetic jets, enabling precise control over the flow dynamics. To present a clear understanding of the DRL-based AFC framework, we provide a detailed explanation of each key component: the agent, the environment, state observations, action selection, and the reward mechanism. By elaborating on these elements, we aim to clarify their roles within the system and illustrate how they interact. This detailed breakdown enhances our ability to analyze and understand the application of DRL in optimizing AFC performance.

### v5

@@ 2.2 Circular cylinder environment
Figure 2: For a pair of jets, five arrangements are considered: (a) $J_{1}^{1}$ and $J_{1}^{2}$, (b) $J_{2}^{1}$ and $J_{2}^{2}$, (c) $J_{3}^{1}$ and $J_{3}^{2}$, (d) $J_{4}^{1}$ and $J_{4}^{2}$, (e) $J_{5}^{1}$ and $J_{5}^{2}$. For two pairs of jets, two arrangements are examined: (f) $J_{3}^{1}$, $J_{3}^{2}$ and $J_{4}^{1}$, $J_{4}^{2}$; (g) $J_{3}^{1}$, $J_{3}^{2}$ and $J_{2}^{1}$, $J_{2}^{2}$. For three pairs of jets, one arrangement is studied: (h) $J_{2}^{1}$, $J_{2}^{2}$, $J_{3}^{1}$, $J_{3}^{2}$, and $J_{4}^{1}$, $J_{4}^{2}$.
In the numerical simulation, the synthetic jets are modeled with a parabolic velocity distribution, where the velocity peaks at the jet center and decreases to zero at the edges. Each jet spans an angular width of $\omega=10^{\circ}$ and is oriented perpendicularly to the cylinder surface. The control action directly adjusts the amplitude of this parabolic profile, thereby regulating the jet intensity. A positive value of the control action represents that the synthetic jet is in a blowing state, and a negative value of the control action represents that the synthetic jet is in a suction state. To maintain global mass conservation, the net mass flow rate of the jet pair is constrained such that $V_{\Gamma_{1}}=-V_{\Gamma_{2}}$, ensuring that no net mass is introduced or removed from the system. Through this mechanism, synthetic jet pairs dynamically influence the near-wall flow characteristics, providing an effective means to modulate vortex shedding, reduce drag, and enhance wake stability.
To systematically investigate the influence of synthetic jet positioning on flow control around a circular cylinder, experiments are conducted using three configurations: a single-pair, two-pairs and three-pairs jet arrangement. The specific placements of the jets in each configuration are detailed in figure 2. Through this mechanism, synthetic jet pairs dynamically influence the near-wall flow characteristics, providing an effective means to modulate vortex shedding, reduce drag, and enhance wake stability.
  • 1.
    Single-Pair Configuration: Five distinct jet placement schemes are evaluated, where the jet pair is denoted as $J_{1}^{1}$ and $J_{1}^{2}$. The subscript represents the pair index, while the superscript distinguishes individual jets within the pair. The jets are positioned at azimuthal angles of $60^{\circ}$, $75^{\circ}$, $90^{\circ}$, $105^{\circ}$, and $120^{\circ}$, enabling an assessment of the impact of individual jet placement on flow control effectiveness.
  • 2.
    Two-pairs Configuration: Two placement schemes are considered to analyze the interaction effects between jet pairs. In this setup, additional jets are introduced to examine how multiple jets influence vortex dynamics and flow stabilization compared to the single-pair case.
  • 3.
    Three-Pairs Configuration: This configuration activates six jets, designated as $J_{2}^{1}$, $J_{2}^{2}$, $J_{3}^{1}$, $J_{3}^{2}$, $J_{4}^{1}$, and $J_{4}^{2}$. The objective is to evaluate the cumulative effect of multiple jet pairs on flow control and identify potential synergies that enhance performance.
@@ 2.3 Square cylinder environment
Figure 3: Schematic of the computational domain and boundary conditions. (a) Establishment of the coordinate system and detailed dimensions of the computational domain. (b) Placement of the synthetic jets near the trailing corner point. (c) The mesh generation scheme around the square object, illustrating the distribution of high-density mesh in regions prone to flow separation, and (d) A close-up view of the mesh structure.
The computational domain for simulating the flow around a square cylinder is defined as a rectangle with dimensions $40D\times 20D$, where $D$ denotes the side length of the square cylinder. A Cartesian coordinate system is established with the center of the square cylinder as the origin, and the positive $x$-axis represents the streamwise direction. The domain extends $10D$ upstream and $30D$ downstream from the cylinder center, while the vertical extent spans ±$\pm 10D$. As illustrated in figure 3(a), the computational domain and boundary conditions are configured to minimize boundary effects and accurately capture wake dynamics. A uniform velocity profile is imposed at the inlet boundary $\Gamma_{\text{in}}$, corresponding to a Reynolds number of 100. The velocity of the synthetic jets is directed normal to the cylinder walls, as depicted in figure 3(b). A no-slip boundary condition is applied to the remaining surfaces of the square cylinder. At the outlet boundary $\Gamma_{\text{out}}$, a Neumann condition is enforced to maintain a zero-stress vector, ensuring a physically consistent outflow. Additionally, far-field conditions are imposed on the upper and lower boundaries to minimize their influence on the flow around the cylinder.
To ensure numerical stability, a time step of $\Delta t=0.0005$ is selected based on extensive validation, striking a balance between temporal accuracy and compliance with the CFL condition. This choice minimizes numerical errors and enhances the fidelity of the simulation. The grid independence study for the numerical simulation of flow around the square cylinder is presented in Appendix 7. The computational mesh used in subsequent simulations, shown in figure 3(c), gradually coarsens radially outward from the square cylinder to optimize computational efficiency. Meanwhile, figure 3(d) highlights local grid refinement near the cylinder.
@@
To investigate the sensitivity of control performance to synthetic jet placement and the synergistic effects of multiple jet pairs, we design configurations where one, two, or three pairs of synthetic jets are activated simultaneously. For the single-pair jet configuration, four distinct placement schemes are considered. Each jet is named according to its position, with $J_{1}^{1}$ and $J_{1}^{2}$ representing jets located near the front corners. Additionally, configurations are designed where two pairs and three pairs of jets operate simultaneously. In multi-pair jet configurations, the DRL algorithm assigns independent control actions to each jet pair, allowing for a detailed evaluation of how different placements affect flow control effectiveness. The various jet placement schemes are illustrated in figure 4.
@@ 2.4 Deep Reinforcement Learning
Deep learning, a branch of machine learning, utilizes deep neural networks with input, hidden, and output layers to process raw data, extract features, and generate task-specific predictions LeCun et al. (2015); Granter et al. (2017). Reinforcement learning complements this by enabling agents to learn optimal actions in dynamic environments through trial and error, refining strategies by maximizing cumulative rewards Mnih (2013); Wang et al. (2020).
Deep learning is a branch of machine learning that uses deep neural networks with input, hidden, and output layers LeCun et al. (2015). These networks process raw data to extract relevant features and generate predictions for specific tasks Granter et al. (2017). Reinforcement Learning is a machine learning paradigm where an agent learns optimal actions in a dynamic environment through trial and error Mnih (2013). By maximizing cumulative rewards via feedback, the agent iteratively refines its strategy to adapt to environmental complexities and make effective decisions Wang et al. (2020).
2.4.1 Foundation
As illustrated in figure 5(a), the agent interacts with the environment by taking actions ($a\in\mathbb{A}$) based on its policy ($\pi(a|s)$), which guides action selection from the current state ($s\in\mathbb{S}$) François-Lavet et al. (2018). The environment provides feedback as rewards, enabling the agent to optimize its strategy. The state value function ($V(s)$) estimates the expected cumulative reward from a state, while the action value function ($Q(s,a)$) evaluates the expected reward from a specific action in a given state Wang et al. (2020).
2.4.2 Mathematical model
RL relies on a finite Markov Decision Process (MDP), defined by the tuple $(S,A,P,R,\gamma)$, where $S$ is the state space, $A$ the action space, $P$ the state transition probability, $R$ the reward function, and $\gamma$ the discount factor Puterman (1990); Altman (1999). At each time step $i$, the agent observes state $s_{i}$, selects action $a_{i}$, transitions to state $s_{i+1}$ with probability $p(s_{i+1}|s_{i},a_{i})$, and receives reward $R(s_{i},a_{i},s_{i+1})$. Decision-making is guided by a policy $\pi(a_{i}|s_{i})$, defining the probability of choosing $a_{i}$ in state $s_{i}$.
2.4.3 Classification
@@
In reinforcement learning, the value function $V(s)$ and the action-value function $Q(s,a)$ serve as critical metrics for assessing the potential of states and state-action pairs to yield high cumulative rewards. These are defined as follows Szepesvári and Littman (1999):
[…]
@@
where $\mathbb{E}_{\pi}$ denotes the expectation under the policy $\pi$, $\gamma$ is the discount factor ($0\leq\gamma\leq 1$), and $r_{t}$ is the immediate reward at time step $t$. $s_{0}=s$ indicates that the initial state is $s$, and $a_{0}=a$ specifies the initial action for $Q(s,a)$. These functions collectively describe the expected cumulative rewards for states and state-action pairs, guiding the agent in optimizing its policy to maximize long-term returns.
RL is broadly categorized into value-based and policy-based methods. Value-based methods, such as Q-learning and DQN, learn $V(s)$ or $Q(s,a)$ to estimate cumulative rewards and select optimal actions by maximizing these values Mnih et al. (2015); Altman (1999). Policy-based methods directly optimize a policy function ($\pi(a|s)$) to maximize rewards, excelling in continuous action spaces and stochastic policies. Examples include policy gradient methods, TRPO, and PPO Silver et al. (2014); Schulman et al. (2015, 2017). The structure of policy-based training is shown in figure 5(b).
Reinforcement learning is broadly categorized into value-based and policy-based methods. Value-based methods, such as Q-learning and Deep Q-Networks, learn value functions like $V(s)$ or $Q(s,a)$ to estimate cumulative rewards. Optimal actions are selected by maximizing these values, making value-based approaches particularly effective in discrete action spaces Mnih et al. (2015); Altman (1999). Policy-based methods take a different approach by directly optimizing a policy function $\pi(a|s)$ to maximize expected rewards. These methods are well-suited for problems involving continuous action spaces and stochastic policies. Policy gradient methods, including Trust Region Policy Optimization and Proximal Policy Optimization, refine policies through iterative updates, ensuring stability and efficiency in training Silver et al. (2014); Schulman et al. (2015, 2017). Figure 5(b) illustrates the structure of policy-based training, highlighting its iterative nature and optimization process.
2.4.4 policy optimization algorithm
PPO is a robust policy optimization algorithm that stabilizes performance by constraining policy update amplitudes. It uses a clipped objective to keep the probability ratio of the current and old policies within a safe range, preventing performance fluctuations. By leveraging the advantage function and optimizing via stochastic gradient descent, PPO efficiently learns the optimal policy Schulman et al. (2017).
$L^{\text{CLIP}}(\theta)=\hat{\mathbb{E}}_{t}\left[\min\left(r_{t}(\theta)\hat{A}_{t},\text{clip}(r_{t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{t}\right)\right],$ (10)
where $r_{t}(\theta)=\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}$ measures policy change, and the advantage estimate $\hat{A}_{t}$ evaluates the benefit of action $a_{t}$ in state $s_{t}$. A clipping mechanism, $\text{clip}(r_{t}(\theta),1-\epsilon,1+\epsilon)$, ensures stability by restricting policy updates within $[1-\epsilon,1+\epsilon]$, where $\epsilon$ controls the update range.
2.5 Enhanced Active Flow Control
Through a DRL-based AFC framework, we integrate a CFD environment with DRL algorithms to address flow control problems. The PPO agent is employed to train a control policy that optimizes real-time actuation of two synthetic jets positioned on the cylinder’s surface, with the objective of minimizing fluid forces. The agent perceives the CFD environment through strategically placed flow probes that measure velocity and pressure. Based on these observations, the agent dynamically adjusts the mass flow rate of the synthetic jets to manipulate the flow field. This framework, comprising the agent, environment, state observations, action selection, and reward mechanism, is meticulously detailed to elucidate the intricate interplay among these components in enhancing the performance of AFC.